OpenStudy (anonymous):

A construction company was contracted to build a cylindrical storage tank with a diameter of 60 feet and height of 16 feet. Just before construction was to begin, it was learned that a new law allows tanks to be a maximum of 9 feet in height. If the tank is redesigned to meet the law, but has the same volume, what is its new diameter? a.)106 feet, 8 inches b.)80 feet c.)40 feet d.)53 feet, 4 inches

OpenStudy (radar):

V=pi r^2 h, v=pi 30^2 16 v= pi r^2 9 v=v p r^2 9 = pi 30^2 16 Divide thru by pi r^2 9 =30^2 16 r^2 =(30^2 16)/9=1600 r=sq rt 1600 r=_______

OpenStudy (anonymous):

I got 40 feet. Is that correct?

OpenStudy (radar):

Compute r and double to get diameter as the question is asking for diameter.

OpenStudy (anonymous):

Oh okay.

OpenStudy (radar):

The radius is indeed 40 feet

OpenStudy (radar):

Compute the diameter.

OpenStudy (anonymous):

So it's 80 for the diameter!

OpenStudy (radar):

Your calculations and reasoning is good.

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (radar):

You're welcome hope you followed with understanding, did you see how we didn't have to mess with the irrational pi?

OpenStudy (anonymous):

Yes ahaha

OpenStudy (radar):

Good luck in your studies.

OpenStudy (anonymous):

Thanks